OpenStudy (zale101):

Use any method to find the relative extrema of f(x)=2x+3x^(2/3)

OpenStudy (anonymous):

\[2+\frac{ 2 }{ x ^{\frac{ 1 }{ 3 }} }=0,\frac{ 2 }{ x ^{\frac{ 1 }{ 3 }} }=-2,x ^{\frac{ 1 }{ 3 }}=\frac{ 2 }{ -2 }=-1\] cubing x=-1 \[\frac{ d^2y }{ dx^2 }=?~at~x=-1\] ?

OpenStudy (zale101):

Thanks!

OpenStudy (zale101):

So, when finding the relative extrema, you use the critical points for the first derivative and test it to the second derivative? :)

OpenStudy (anonymous):

correct.

OpenStudy (zale101):

Thanks again.

OpenStudy (anonymous):

yw